OpenStudy (anonymous):

Find the x & y intercept 8.4x-33.6y=100.8

OpenStudy (anonymous):

put x=0 to get y-intercept put y=0 to get x-intercept

OpenStudy (anonymous):

@NatalieLove can you do this?

OpenStudy (anonymous):

I don't have a calculator

OpenStudy (anonymous):

you can use the calculator in your computer

OpenStudy (anonymous):

I'm in my phone

OpenStudy (anonymous):

ok can i do the caalculation

OpenStudy (anonymous):

@NatalieLove u hav the calculator on computer just type calculator in run

OpenStudy (anonymous):

@akash_809 she is using in phone

OpenStudy (anonymous):

@uzumakhi dont phone have clculators...i bet every phone has that

OpenStudy (anonymous):

dont tell me..........tell to her.........lol

OpenStudy (anonymous):

What ever ill probably fail the test

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Nothing much

OpenStudy (anonymous):

you have 3 numbers to play with; 2 angles and a side tan(30) = db/bc ; we know bc = 12 tan(30) = db/12 solve for db 12 tan(30) = db ----------------------- tan(30+15) = ab/bc tan(45) = ab/12 12 tan(45) = ab ----------------------- ab - db = ad 12 tan(45) - 12tan(30) = ad